# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ Acknowledgments
In this section, we detail how the circuit parameters are estimated by analyzing the dependency of each mode resonance frequency against DC bias fluxes, in the neighborhood of the ATS operating points (see Fig. 2a, operating points appearing as saddle points in the color map). Our model Hamiltonian is derived from the three-mode effective circuit pictured in Fig. 2a (obtained after Foster synthesis of the environment seen by the SQUID) and includes the non-idealities listed in Sec. S2, it reads
@@
$\begin{split}{\bf H}=\sum_{i}\hbar\omega_{i}{\bf a}_{i}^{\dagger}{\bf a}_{i}&-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\bm{\varphi}-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\bm{\varphi}-\varphi_{\Delta})\\ &+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\bm{\varphi}-2\varphi_{\Delta})+\frac{E_{L}}{N^{2}}\big{(}\cos(\bm{\varphi})+\frac{1}{2}\bm{\varphi}^{2}\big{)}\end{split}$ (S19)
@@
In this expression, $i\in\{a,b,r\}$ labels each mode, $\bm{\varphi}=\sum_{i}\varphi_{i}^{ZPF}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ is the reduced phase across the SQUID, $\varphi_{i}^{ZPF}=(\pi Z_{i}/R_{q})^{1/2}$ represents the vacuum fluctuations of the phase of each mode across the SQUID. We have defined the reduced common and differential bias fluxes $(\varphi_{\Sigma},\varphi_{\Delta})$ in Eq. (S7). As detailed in Sec. S2, we have allowed for the ATS junctions to be imbalanced with respective energies $E_{J_{1}}$ and $E_{J_{2}}$ and defined $E_{\Sigma}=E_{J_{1}}+E_{J_{2}}$ and $E_{\Delta}=E_{J_{1}}-E_{J_{2}}$. The last term on the first line stems from this imbalance (see Sec. S2). The first term on the second line stems from series inductances in the arms of the SQUID. We have assumed that series inductances are identical on both arms (inductive energy $\epsilon_{L}$) and defined $e_{\Sigma}=(E^{2}_{J_{1}}+E^{2}_{J_{2}})/(4\epsilon_{L})$ (a last term in $(E^{2}_{J_{1}}-E^{2}_{J_{2}})/(4\epsilon_{L})$ is neglected). Finally, the last term on the second line stems from the non-linear contribution of the superinductors potential (each a chain of $N=25$ for a total inductance $E_{L}$ [^10] This term is only valid in the RWA since the modes phase fluctuations can have opposed signs accross both inductors. However, the only contributions that survives the RWA do not depend on this sign.
@@
In practice, we first record the frequency of the readout mode over a full period of modulation for both fluxes. After locating three saddle points on the recorded map, we find the linear transform relating DC currents applied to the circuit and reduced fluxes $\varphi_{\Sigma}$ and $\varphi_{\Delta}$. We then measure all three modes resonance frequencies nearby the saddle points. More specifically, the fluxes are varied along cuts intersecting $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,\pm\pi/2)$, and we record the transition frequencies $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ of the memory and the lowest transition frequency of the buffer and readout modes. In order to facilitate the fit of these datasets with the model Hamiltonian (S19), we first fit the measured frequencies along each cut with a parabola, and report the extracted curvature against cut angle in Fig. S12. We subsequently fit the curvature pattern with our model. We first detail a simplified classical analysis of the circuit around the frequency saddle points, which is the rationale for the fitting procedure to the full quantum model detailed in the following section.
@@
We here compute simple analytical formula for the frequency of each mode against externally applied fluxes in the neighborhood of the saddle points, using a classical model. We neglect all circuit non-idealities ($E_{\Delta}=e_{\Sigma}=E_{L}/N^{2}=0$ in (S19)). Moreover, for simplicity, we consider two only modes $a$ and $b$ with phases $\varphi_{a}$ and $\varphi_{b}$. We will generalize the obtained formula to the three-mode case at the end of the calculation.
@@
$V(\varphi_{a},\varphi_{b})=\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}\sin(\epsilon_{\Sigma})\sin(\varphi_{a}+\varphi_{b}-\epsilon_{\Delta})$ (S20)
@@
$\epsilon_{\Sigma}$ $=\epsilon\cos(\theta)$ (S21)
@@
$V(\varphi_{a},\varphi_{b})=$ $(\displaystyle\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}c\epsilon\Big{(}$S22)
@@
$\varphi_{a}^{m}$ $=\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$ (S23)
@@
$\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})=0$ (S24)
@@
Figure S11: Effective circuit corresponding to the inductive potential encoded by the Hessian matrix (S27) and the charging energies found by Foster synthesis of the SQUID environment.
@@
$\varphi_{a}^{m}$ $=-\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})$ (S25)
@@
$\frac{\partial^{2}V}{\partial\varphi_{a}^{2}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}$ $=E_{L_{a}}\big{(}1+\epsilon^{2}r_{a}c(s+r_{a}c+r_{b}c)\big{)}+\mathcal{O}(\epsilon^{3})$ (S26)
@@
$\begin{pmatrix}E_{L_{a}}+F&F\\ F&E_{L_{b}}+F\end{pmatrix}$ (S27)
at second order in $\epsilon$. Since the modes charging energies do not vary as a function of flux, this effective potential combined with the charging energies correspond to the circuit pictured in Fig. S11, where we have defined $L_{F}=\frac{\varphi_{0}^{2}}{F}$. We can write down the circuit Lagrangian and find the two dressed mode frequencies
$\omega_{\pm}^{2}=\frac{\tilde{\omega}_{a}^{2}+\tilde{\omega}_{b}^{2}}{2}\pm\frac{\big{(}(\tilde{\omega}_{a}^{2}-\tilde{\omega}_{b}^{2})^{2}+4\omega_{F}^{4}\big{)}^{1/2}}{2}$ (S28)
@@
$\tilde{\omega}_{a}$ $=\omega_{a}\big{(}1+\frac{\epsilon^{2}}{2}r_{a}c(r_{a}c+r_{b}c+s)\big{)}$ (S29)
@@
$C_{i}(\theta)=\omega_{i}r_{i}\cos(\theta)\big{(}(r_{a}+r_{b}+r_{c})\cos(\theta)+\sin(\theta)\big{)}$ (S30)
@@
  • Imbalance of the SQUID junctions modifies the inductive Hessian matrix (S27) following
    $H=\begin{pmatrix}E_{L_{a}}-E_{\Delta}+F&-E_{\Delta}+F\\ -E_{\Delta}+F&E_{L_{b}}-E_{\Delta}+F\end{pmatrix}$ (S31)
@@
Figure S12: Curvature of the frequency maps around the saddle points a-b) Measured (filled circles) and simulated (plain lines) curvature of the memory and buffer low-energy transition frequency at the two non-equivalent saddle points, against cut angle $\theta$ (defined in Eq. (S21)). Simulated values use the model Hamiltonian (S19) fitted to the data. Dashed lines represent the predictions from the classical model (S30) with the same parameters. These only reproduce qualitatively the data. Note that the curvature of the readout mode has been enhanced by a factor 20 for visibility. c) Predictions from the quantum and classical model for a hypothetical circuit whose modes have five times lower phase fluctuations. Here, the two models agree quantitatively.
In the classical model detailed above, we have neglected the vacuum fluctuations of each mode. Therefore, we do not expect the predictions to be accurate when the modes vacuum fluctuations approach 1, which is the case in our experiment. After a first coarse estimate of the circuit parameters using our classical model, we refine our analysis by numerically diagonalizing the Hamiltonian (S19). In Fig. S12a-b, we represent the curvature $C_{i}$ of each mode resonance frequency around the saddle point against cut angle $\theta$ (see Eq. (S21)), together with the result of these simulations (for the estimated circuit parameters) and the predictions from the classical analysis. The latter only qualitatively reproduce the data. For comparison, we represent in Fig. S12c the predictions from the Hamiltonian diagonalization against classical predictions for a circuit whose modes have five times smaller vacuum fluctuations across the SQUID. Here, the agreement is quantitative.
@@
S4.2 Dynamics under the four-to one photon exchange interaction
@@
We start from the circuit Hamiltonian (S19), which is time-dependent for flux biases $(\varphi_{\Sigma},\varphi_{\Delta})=(-\frac{\pi}{2}+\epsilon(t),-\frac{\pi}{2})$ that include a small AC component ($\epsilon(t)\ll 1$). Neglecting the readout mode, expanding at first order in $\epsilon$ and including a direct linear drive on the buffer (needed to reproduce the buffer reflection spectra in Fig. 3c), we get
$\begin{split}{\bf H}(t)=&\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b^{\dagger}}{\bf b}+\epsilon(t)E_{\Sigma}\sin(\bm{\varphi})\\ &+E_{\Delta}\cos(\bm{\varphi})+e_{\Sigma}\cos(2\bm{\varphi})+\frac{E_{L}}{N^{2}}\big{(}\cos(\bm{\varphi})+\frac{1}{2}\bm{\varphi}^{2}\big{)}\\ &+\epsilon_{b}(t)({\bf b}+{\bf b^{\dagger}})\end{split}$ (S32)
where $\bm{\varphi}=\varphi_{a}^{ZPF}({\bf a}+{\bf a^{\dagger}})+\varphi_{b}^{ZPF}({\bf b}+{\bf b^{\dagger}})$. Our goal is to derive an effective static Hamiltonian allowing for numerical simulation. To this end, we move to a rotating frame for both modes, and select resonant or near-resonant terms.
We assume the pump to be near-resonant with the four-to-one photon transition $\epsilon(t)=\xi e^{i(4\tilde{\omega}_{a}-\tilde{\omega}_{b}+\Delta)t}+\mathrm{c.c}.$ and the linear drive to be near resonant with the buffer $\epsilon_{b}(t)=\xi_{b}e^{i(\tilde{\omega}_{b}+\delta)t}+\mathrm{c.c}.$ with $\Delta,\delta\ll\tilde{\omega}_{a},\tilde{\omega}_{b}$. Here, $\tilde{\omega}_{a,b}$ denote the modes resonance frequencies dressed by the pump, which are assumed to be perfectly calibrated by the four-step procedure detailed in Sec. S3.2. When simulating the time-decay of the state $|4\rangle_{a}$, we thus set $\Delta=0$ and $\xi_{b}=0$ (pump on resonance, no buffer drive, see Fig. 3b). When reproducing the buffer reflection spectra (Fig. 3c), $\delta$ is the detuning between the applied buffer drive and the buffer resonance frequency dressed by the pump. We give a non-zero value to $\Delta=\omega_{b_{3}}-\omega_{b_{1}}$ to account for the slightly off-resonant pump (we use the convention of the four-step calibration procedure and recall that the spectra are recorded at step 3). In practice, this small detuning has no visible impact on the simulation results.
@@
$\begin{split}a&\rightarrow a^{-i\big{(}\tilde{\omega}_{a}+\frac{\Delta+\delta}{4}\big{)}t}\\ b&\rightarrow b^{-i(\tilde{\omega}_{b}+\delta)t}.\end{split}$ (S33)
First considering the terms on the second line of the Hamiltonian (S32) stemming from non-idealities of the circuit, we discard contributions that are fast-oscillating in the rotating frame, i.e. which do not preserve photon number. This amounts to keeping only diagonal terms in the Fock basis and we end up with a diagonal Hamiltonian
${\bf H}_{\chi}=\sum_{n,m}\chi_{nm}|n\rangle_{a}|m\rangle_{b}\langle n|_{a}\langle m|_{b}$ (S34)
where the dispersive shifts $\chi_{nm}$ are computed from the non-driven version of Hamiltonian (S32) (setting $\epsilon=\epsilon_{b}=0$). In detail, we first diagonalize the non-driven Hamiltonian and label $(n,m)$ each eigenstate by identifying it to the Fock state $|n\rangle_{a}|m\rangle_{b}$ it maximally overlaps with  [^11] The maximum overlap (defined for two states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ as $|\langle\psi_{1}|\psi_{2}\rangle|^{2}$) is larger than 0.9995 for all $(n,m)$. Next, we label each state energy $E_{n,m}$ after offsetting the energy scale such that $E_{0,0}=0$. Finally, we find the dispersive shift $\chi_{nm}=E_{n,m}-n\frac{E_{4,0}}{4}-mE_{0,1}$. We note that with our definitions, the dispersive Hamiltonian ${\bf H}_{\chi}$ does not contribute to the energy splitting between $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$, which is expected given our choice of rotating frames. We also note that by computing ${\bf H}_{\chi}$ from the non-driven Hamiltonian (S32), we have neglected a small renormalization by the pump. In practice, the exact value of ${\bf H}_{\chi}$ does not significantly impact numerical simulations in the regime we are in, as detailed below.
Now turning to the term $\epsilon(t)E_{\Sigma}\sin(\bm{\varphi})$, we get in the RWA a resonant four-to-one photon exchange interaction at rate $g_{4}=\frac{E_{\Sigma}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ (see Sec. S1.1). Overall, we get the effective static Hamiltonian
${\bf H}_{RWA}=-\frac{\Delta+\delta}{4}{\bf a}^{\dagger}{\bf a}-\delta{\bf b}^{\dagger}{\bf b}+{\bf H}_{\chi}+g_{4}({\bf a}^{\dagger 4}{\bf b}+{\bf a}^{4}{\bf b^{\dagger}})+\xi_{b}{\bf b}+\xi_{b}^{\ast}{\bf b^{\dagger}}$ (S35)
@@
  • $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$. These dissipators account for the relaxation of population in state $|k\rangle$ through single-photon emission. The rates $\gamma_{k}$ for $k\leq 3$ are independently calibrated for each pump amplitude (see Sec. S3.3). Since we do not consider any other decay mechanism from these states, they are identical to the rates $\Gamma_{k}$ reported in Fig. 4. Note that their values do not follow the expected linear scaling in $k$, preventing us from modeling single-photon dissipation with a single dissipator $\mathcal{D}[{\bf a}]$. Given the memory anharmonicity, this behavior is expected if the coupling to the bath in which single-photons are emitted—or the bath density of states—is not constant in frequency. As for $\gamma_{4}$, it cannot be unambiguously estimated for a non-zero pump amplitude since decay from state $|4\rangle_{a}$ under pump is dominated by the engineered four-photon dissipation channel. We therefore assume $\gamma_{4}$ to be constant over all considered pump amplitudes.
@@
  • $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$. This dissipator accounts for pure dephasing of the memory. The dephasing rate is fitted together with the datasets presented in Fig. 3 and is in good agreement with an independent measurement of the dephasing of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition. Note that this rate is enhanced by the application of the off-resonant pump employed to shift the memory transitions away from spurious TLS resonances (see Sec. S5) and was seen to fluctuate over the course of hours.
@@
  • A scaling factor accounting for the cryostat input and output lines attenuation (determining the amplitude of off-resonant waves reflected from the buffer in Fig. 3c and Fig. S10).

### v3

@@ D.1 Frequency maps
In this section, we detail how the circuit parameters are estimated by analyzing the dependency of each mode resonance frequency against DC bias fluxes, in the neighborhood of the ATS operating points (see Fig. 2a, operating points appearing as saddle points in the color map). Our model Hamiltonian is derived from the three-mode effective circuit pictured in Fig. 2a (obtained after Foster synthesis of the environment seen by the SQUID) and includes the non-idealities listed in Sec. B, it reads
@@
$\begin{split}{\bf H}=&\sum_{i}\hbar\omega_{i}{\bf a}_{i}^{\dagger}{\bf a}_{i}\\ &-E_{\Sigma}\cos(\varphi_{\Sigma})\cos(\boldsymbol{\varphi}-\varphi_{\Delta})-E_{\Delta}\sin(\varphi_{\Sigma})\sin(\boldsymbol{\varphi}-\varphi_{\Delta})\\ &+e_{\Sigma}\cos(2\varphi_{\Sigma})\cos(2\boldsymbol{\varphi}-2\varphi_{\Delta})+\frac{E_{L}}{N^{2}}\big(\cos(\boldsymbol{\varphi})+\frac{1}{2}\boldsymbol{\varphi}^{2}\big)\end{split}$ (30)
@@
In this expression, $i\in\{a,b,r\}$ labels each mode, $\boldsymbol{\varphi}=\sum_{i}\varphi_{i}^{ZPF}({\bf a}_{i}+{\bf a}_{i}^{\dagger})$ is the reduced phase across the SQUID, $\varphi_{i}^{ZPF}=(\pi Z_{i}/R_{q})^{1/2}$ represents the vacuum fluctuations of the phase of each mode across the SQUID. We have defined the reduced common and differential bias fluxes $(\varphi_{\Sigma},\varphi_{\Delta})$ in Eq. (12). As detailed in Sec. B, we have allowed for the ATS junctions to be imbalanced with respective energies $E_{J_{1}}$ and $E_{J_{2}}$ and defined $E_{\Sigma}=E_{J_{1}}+E_{J_{2}}$ and $E_{\Delta}=E_{J_{1}}-E_{J_{2}}$. The last term on the first line stems from this imbalance (see Sec. B). The first term on the second line stems from series inductances in the arms of the SQUID. We have assumed that series inductances are identical on both arms (inductive energy $\epsilon_{L}$) and defined $e_{\Sigma}=(E^{2}_{J_{1}}+E^{2}_{J_{2}})/(4\epsilon_{L})$ (a last term in $(E^{2}_{J_{1}}-E^{2}_{J_{2}})/(4\epsilon_{L})$ is neglected). Finally, the last term on the second line stems from the non-linear contribution of the superinductors potential (each a chain of $N=25$ for a total inductance $E_{L}$[75].
@@
In practice, we first record the frequency of the readout mode over a full period of modulation for both fluxes. After locating three saddle points on the recorded map, we find the linear transform relating DC currents applied to the circuit and reduced fluxes $\varphi_{\Sigma}$ and $\varphi_{\Delta}$. We then measure all three modes resonance frequencies nearby the saddle points. More specifically, the fluxes are varied along cuts intersecting $(\varphi_{\Sigma},\varphi_{\Delta})=(-\pi/2,\pm\pi/2)$, and we record the transition frequencies $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ of the memory and the lowest transition frequency of the buffer and readout modes. In order to facilitate the fit of these datasets with the model Hamiltonian (30), we first fit the measured frequencies along each cut with a parabola, and report the extracted curvature against cut angle in Fig. 16. We subsequently fit the curvature pattern with our model. We first detail a simplified classical analysis of the circuit around the frequency saddle points, which is the rationale for the fitting procedure to the full quantum model detailed in the following section.
@@
We here compute simple analytical formula for the frequency of each mode against externally applied fluxes in the neighborhood of the saddle points, using a classical model. We neglect all circuit non-idealities ($E_{\Delta}=e_{\Sigma}=E_{L}/N^{2}=0$ in (30)). Moreover, for simplicity, we consider two only modes $a$ and $b$ with phases $\varphi_{a}$ and $\varphi_{b}$. We will generalize the obtained formula to the three-mode case at the end of the calculation.
@@
$V(\varphi_{a},\varphi_{b})=\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}\sin(\epsilon_{\Sigma})\sin(\varphi_{a}+\varphi_{b}-\epsilon_{\Delta})$ (31)
@@
$\epsilon_{\Sigma}$ $=\epsilon\cos(\theta)$ (32)
@@
$V(\varphi_{a},\varphi_{b})=$ $\frac{1}{2}E_{L_{a}}\varphi_{a}^{2}+\frac{1}{2}E_{L_{b}}\varphi_{b}^{2}+E_{\Sigma}c\epsilon\Big($ (33)
@@
$\varphi_{a}^{m}$ $=\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$ (34)
@@
$\alpha_{a}\epsilon+\beta_{a}\epsilon^{2}+\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})=0$ (35)
@@
Figure 15: Effective circuit corresponding to the inductive potential encoded by the Hessian matrix (38) and the charging energies found by Foster synthesis of the SQUID environment.
@@
$\varphi_{a}^{m}$ $=-\epsilon r_{a}c+\mathcal{O}(\epsilon^{3})$ (36)
@@
$\frac{\partial^{2}V}{\partial\varphi_{a}^{2}}|_{(\varphi_{a}^{m},\varphi_{b}^{m})}$ $=E_{L_{a}}\big(1+\epsilon^{2}r_{a}c(s+r_{a}c+r_{b}c)\big)+\mathcal{O}(\epsilon^{3})$ (37)
@@
$\begin{pmatrix}E_{L_{a}}+F&F\\ F&E_{L_{b}}+F\end{pmatrix}$ (38)
at second order in $\epsilon$. Since the modes charging energies do not vary as a function of flux, this effective potential combined with the charging energies correspond to the circuit pictured in Fig. 15, where we have defined $L_{F}=\frac{\varphi_{0}^{2}}{F}$. We can write down the circuit Lagrangian and find the two dressed mode frequencies
$\omega_{\pm}^{2}=\frac{\tilde{\omega}_{a}^{2}+\tilde{\omega}_{b}^{2}}{2}\pm\frac{\big((\tilde{\omega}_{a}^{2}-\tilde{\omega}_{b}^{2})^{2}+4\omega_{F}^{4}\big)^{1/2}}{2}$ (39)
@@
$\tilde{\omega}_{a}$ $=\omega_{a}\big(1+\frac{\epsilon^{2}}{2}r_{a}c(r_{a}c+r_{b}c+s)\big)$ (40)
@@
$C_{i}(\theta)=\omega_{i}r_{i}\cos(\theta)\big((r_{a}+r_{b}+r_{c})\cos(\theta)+\sin(\theta)\big)$ (41)
@@
  • Imbalance of the SQUID junctions modifies the inductive Hessian matrix (38) following
    $H=\begin{pmatrix}E_{L_{a}}-E_{\Delta}+F&-E_{\Delta}+F\\ -E_{\Delta}+F&E_{L_{b}}-E_{\Delta}+F\end{pmatrix}$ (42)
@@
Figure 16: Curvature of the frequency maps around the saddle points a-b) Measured (filled circles) and simulated (plain lines) curvature of the memory and buffer low-energy transition frequency at the two non-equivalent saddle points, against cut angle $\theta$ (defined in Eq. (32)). Simulated values use the model Hamiltonian (30) fitted to the data. Dashed lines represent the predictions from the classical model (41) with the same parameters. These only reproduce qualitatively the data. Note that the curvature of the readout mode has been enhanced by a factor 20 for visibility. c) Predictions from the quantum and classical model for a hypothetical circuit whose modes have five times lower phase fluctuations. Here, the two models agree quantitatively.
In the classical model detailed above, we have neglected the vacuum fluctuations of each mode. Therefore, we do not expect the predictions to be accurate when the modes vacuum fluctuations approach 1, which is the case in our experiment. After a first coarse estimate of the circuit parameters using our classical model, we refine our analysis by numerically diagonalizing the Hamiltonian (30). In Fig. 16a-b, we represent the curvature $C_{i}$ of each mode resonance frequency around the saddle point against cut angle $\theta$ (see Eq. (32)), together with the result of these simulations (for the estimated circuit parameters) and the predictions from the classical analysis. The latter only qualitatively reproduce the data. For comparison, we represent in Fig. 16c the predictions from the Hamiltonian diagonalization against classical predictions for a circuit whose modes have five times smaller vacuum fluctuations across the SQUID. Here, the agreement is quantitative.
@@
D.2 Dynamics under the four-to one photon exchange interaction
@@
We start from the circuit Hamiltonian (30), which is time-dependent for flux biases $(\varphi_{\Sigma},\varphi_{\Delta})=(-\frac{\pi}{2}+\epsilon(t),-\frac{\pi}{2})$ that include a small AC component ($\epsilon(t)\ll 1$). Neglecting the readout mode, expanding at first order in $\epsilon$ and including a direct linear drive on the buffer (needed to reproduce the buffer reflection spectra in Fig. 3c), we get
$\begin{split}{\bf H}(t)=&\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b^{\dagger}}{\bf b}+\epsilon(t)E_{\Sigma}\sin(\boldsymbol{\varphi})\\ &+E_{\Delta}\cos(\boldsymbol{\varphi})+e_{\Sigma}\cos(2\boldsymbol{\varphi})+\frac{E_{L}}{N^{2}}\big(\cos(\boldsymbol{\varphi})+\frac{1}{2}\boldsymbol{\varphi}^{2}\big)\\ &+\epsilon_{b}(t)({\bf b}+{\bf b^{\dagger}})\end{split}$ (43)
where $\boldsymbol{\varphi}=\varphi_{a}^{ZPF}({\bf a}+{\bf a^{\dagger}})+\varphi_{b}^{ZPF}({\bf b}+{\bf b^{\dagger}})$. Our goal is to derive an effective static Hamiltonian allowing for numerical simulation. To this end, we move to a rotating frame for both modes, and select resonant or near-resonant terms.
We assume the pump to be near-resonant with the four-to-one photon transition $\epsilon(t)=\xi e^{i(4\tilde{\omega}_{a}-\tilde{\omega}_{b}+\Delta)t}+\mathrm{c.c}.$ and the linear drive to be near resonant with the buffer $\epsilon_{b}(t)=\xi_{b}e^{i(\tilde{\omega}_{b}+\delta)t}+\mathrm{c.c}.$ with $\Delta,\delta\ll\tilde{\omega}_{a},\tilde{\omega}_{b}$. Here, $\tilde{\omega}_{a,b}$ denote the modes resonance frequencies dressed by the pump, which are assumed to be perfectly calibrated by the four-step procedure detailed in Sec. C.2. When simulating the time-decay of the state $|4\rangle_{a}$, we thus set $\Delta=0$ and $\xi_{b}=0$ (pump on resonance, no buffer drive, see Fig. 3b). When reproducing the buffer reflection spectra (Fig. 3c), $\delta$ is the detuning between the applied buffer drive and the buffer resonance frequency dressed by the pump. We give a non-zero value to $\Delta=\omega_{b_{3}}-\omega_{b_{1}}$ to account for the slightly off-resonant pump (we use the convention of the four-step calibration procedure and recall that the spectra are recorded at step 3). In practice, this small detuning has no visible impact on the simulation results.
@@
$\begin{split}a&\rightarrow a^{-i\big(\tilde{\omega}_{a}+\frac{\Delta+\delta}{4}\big)t}\\ b&\rightarrow b^{-i(\tilde{\omega}_{b}+\delta)t}.\end{split}$ (44)
First considering the terms on the second line of the Hamiltonian (43) stemming from non-idealities of the circuit, we discard contributions that are fast-oscillating in the rotating frame, i.e. which do not preserve photon number. This amounts to keeping only diagonal terms in the Fock basis and we end up with a diagonal Hamiltonian
${\bf H}_{\chi}=\sum_{n,m}\chi_{nm}|n\rangle_{a}|m\rangle_{b}\langle n|_{a}\langle m|_{b}$ (45)
where the dispersive shifts $\chi_{nm}$ are computed from the non-driven version of Hamiltonian (43) (setting $\epsilon=\epsilon_{b}=0$). In detail, we first diagonalize the non-driven Hamiltonian and label $(n,m)$ each eigenstate by identifying it to the Fock state $|n\rangle_{a}|m\rangle_{b}$ it maximally overlaps with [71]. Next, we label each state energy $E_{n,m}$ after offsetting the energy scale such that $E_{0,0}=0$. Finally, we find the dispersive shift $\chi_{nm}=E_{n,m}-n\frac{E_{4,0}}{4}-mE_{0,1}$. We note that with our definitions, the dispersive Hamiltonian ${\bf H}_{\chi}$ does not contribute to the energy splitting between $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$, which is expected given our choice of rotating frames. We also note that by computing ${\bf H}_{\chi}$ from the non-driven Hamiltonian (43), we have neglected a small renormalization by the pump. In practice, the exact value of ${\bf H}_{\chi}$ does not significantly impact numerical simulations in the regime we are in, as detailed below.
Now turning to the term $\epsilon(t)E_{\Sigma}\sin(\boldsymbol{\varphi})$, we get in the RWA a resonant four-to-one photon exchange interaction at rate $g_{4}=\frac{E_{\Sigma}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ (see Sec. A.1). Overall, we get the effective static Hamiltonian
$\begin{split}{\bf H}_{RWA}=&-\frac{\Delta+\delta}{4}{\bf a}^{\dagger}{\bf a}-\delta{\bf b}^{\dagger}{\bf b}\\ &+{\bf H}_{\chi}+g_{4}({\bf a}^{\dagger 4}{\bf b}+{\bf a}^{4}{\bf b^{\dagger}})+\xi_{b}{\bf b}+\xi_{b}^{\ast}{\bf b^{\dagger}}\end{split}$ (46)
@@
  • $\gamma_{k}\mathcal{D}[|k-1\rangle\langle k|]$ for $1\leq i\leq 4$. These dissipators account for the relaxation of population in state $|k\rangle$ through single-photon emission. The rates $\gamma_{k}$ for $k\leq 3$ are independently calibrated for each pump amplitude (see Sec. C.3). Since we do not consider any other decay mechanism from these states, they are identical to the rates $\Gamma_{k}$ reported in Fig. 4. Note that their values do not follow the expected linear scaling in $k$, preventing us from modeling single-photon dissipation with a single dissipator $\mathcal{D}[{\bf a}]$. Given the memory anharmonicity, this behavior is expected if the coupling to the bath in which single-photons are emitted—or the bath density of states—is not constant in frequency. As for $\gamma_{4}$, it cannot be unambiguously estimated for a non-zero pump amplitude since decay from state $|4\rangle_{a}$ under pump is dominated by the engineered four-photon dissipation channel. We therefore assume $\gamma_{4}$ to be constant over all considered pump amplitudes.
@@
  • $2\kappa_{\phi}\mathcal{D}[{\bf a^{\dagger}}{\bf a}]$. This dissipator accounts for pure dephasing of the memory. The dephasing rate is fitted together with the datasets presented in Fig. 3 and is in good agreement with an independent measurement of the dephasing of the $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ transition. Note that this rate is enhanced by the application of the off-resonant pump employed to shift the memory transitions away from spurious TLS resonances (see Sec. E) and was seen to fluctuate over the course of hours.
@@
  • A scaling factor accounting for the cryostat input and output lines attenuation (determining the amplitude of off-resonant waves reflected from the buffer in Fig. 3c and Fig. 14).